egiese111304:

You have a credit card with a balance of $1,256.34 and a 9.5% APR. A late payment fee of $29.00 is added to the principal if you pay after 5/28. You make a payment for $200.00 on 5/28. How much interest did you pay?

grezzyy:

u make a payment for $200.00 on 5/28. int is about 1% per mo so is either 9.95 or 8.36 So the equation is (1256.34 * .095/12) so u would have an interest rate of 9.95 1256.34 * .095/12 = 9.95

KyledaGreat:

To calculate the interest paid, we first need to determine the new balance after the payment is made: $1,256.34 - $200.00 = $1,056.34 Next, we can calculate the interest for the period between the payment and the due date (assuming payment is made on the 28th of each month): Interest = (APR / 365) * days * balance In this case: APR = 9.5% Days = 3 (from 5/28 to 5/31) Balance = $1,056.34 Interest = (0.095 / 365) * 3 * $1,056.34 = $0.82 Therefore, the interest paid is $0.82. Adding the $29.00 late payment fee to the principal, the new balance due would be $1,086.16.